OpenStudy (anonymous):

(____)^2=(secx-1)(secx+1)

OpenStudy (anonymous):

what does this also equal

OpenStudy (anonymous):

$$y^2 = (\sec x - 1)(\sec x + 1) = \sec^2 x - 1 = \tan^2 x$$

OpenStudy (anonymous):

so tanx is the answer?

OpenStudy (nincompoop):

no

OpenStudy (anonymous):

hmm...

OpenStudy (loser66):

??? why no???

OpenStudy (anonymous):

im pretty sure it is tanx

OpenStudy (loser66):

I am with you, OK, if you fail, me too. You are not alone, hehehehe

OpenStudy (anonymous):

haha great

OpenStudy (loser66):

Please, be aware of the fact that your credit is real while mine is fake. hehehhe